what is the volume of the cube if the surface area is 36 centimeter cubed?
Answer by nerdybill(7384)

A "cube" means that the height, width, and length are all the same.
Let x = measure of one side of the cube
.
The area of ONE face of the cube is x^2 (x squared).
Their are 6 faces to a cube.
.
The problems said that the total area is 36 sq cm.
6x^2 = 36
x^2 = 6
x = sqrt(6)
.
Volume of the cube is x^3.
x^3 = sqrt(6)*sqrt(6)*sqrt(6)
x^3 = 6*sqrt(6)
x^3 = 14.70 cubic cm